a 13 g sample of an isotope of strontium-90 was formed in 1960 explosion of an atomic bomb at Johnson Island in the pacific test site. in wha year will only 0.625 grams of the strontium-90 remain?


1
Expert's answer
2021-09-22T00:27:02-0400

Radiactive decay law:

"N=N_0e^{-\u03bbt}"

N0 = initial amount of radiactive substances\

N=amount of substance at time t

λ =decay constantly

Half life

"T_{1/2}= \frac{0.693}{λ} \\

N_0=13 \;g \\

N= 0.625 \;g \\

T_{1/2}=28.8 \;years"

Decay constant

"λ = \frac{0.693}{28.8}= 0.024 \;year^{-1} \\

N=N_0e^{-λt} \\

0.625 = 13 \times e^{-0.024t} \\

0.04807 = e^{-0.024t} \\

ln(0.04807) = -0.024t \\

-3.035097 = -0.024t \\

t = \frac{-3.035097}{-0.024} = 126.5 \;years"

1960+126.5 = 2086.5 - year when only 0.625 grams of the strontium-90 will remain